using the equipartition of energy concept estimate the ideal gas Cp(not Cv) in units KB for the low temperature regime where no vibrational modes are activated(if you calculate Cp=42KB, enter 42 as your answer )
1
Expert's answer
2020-06-18T14:51:08-0400

As per equipartition theorem,

The energy of a thermodynamic system in equilibrium is partitioned equally among its degrees of freedom and molar heat capcity is given as

"C_v=\frac{d}{2}R,"

and we know that "C_p=C_v+R"

"C_p=\frac{d}{2}R+R"

where d is the degree of freedom, for ideal gas d=3

Now,

"C_v=\frac{3}{2}R, C_p=\frac{3}{2}R+R=\frac{5}{2}R" , we know that

R=NKB,

"R=6.022\times 10 ^{23}K_B"


"C_P=15\times10^{23}K_B"

so u need to write 15